OpenStudy (anonymous):

Jessica and Cameron have been solving systems of equations with one polynomial function of degree 2 or higher and one linear function. Jessica says there must always be two solutions, and Cameron says there will only be one solution. Using complete sentences, explain how Jessica can be correct, how Cameron can be correct, and how they both can be wrong.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this is like the fifth time i have seen this exact question lets see how they can both be right

OpenStudy (misty1212):

i mean how they can both be wrong

OpenStudy (misty1212):

solve \[y=x^2,y=x-5\] there will be no solution because the line \(y=x-5\) always lies under teh line \(x^2\) here is a picture http://www.wolframalpha.com/input/?i=x%5E2%2Cx-5

OpenStudy (anonymous):

Oooooh I get what's going on, okay. That picture helped. Thanks.

OpenStudy (misty1212):

you want an example where there is only one solution?

OpenStudy (anonymous):

Sure if you'd like to provide one. :)

OpenStudy (misty1212):

if you try \[y=x^2\] and \[y=2x-1\] then if you solve it you get \[x^2=2x-1\\ x^2-2x+1=0\\ (x-1)^2=0\\ x=1\] only one solution here is a picture where you see they touch only at one point http://www.wolframalpha.com/input/?i=x%5E2%2C2x-1

OpenStudy (misty1212):

how about where there is two solutions?

OpenStudy (anonymous):

Sure. Thanks so much, by the way.

OpenStudy (misty1212):

yw

OpenStudy (misty1212):

solve \[y=x^2\\ y=2x+3\] solve and you get \[x^2-2x-3=0\\ (x+1)(x-3)=0\\ x=-1,x=3\] http://www.wolframalpha.com/input/?i=x%5E2%2C2x%2B3

OpenStudy (misty1212):

now you have an example of each, i will let you provide the complete sentences good luck! \[\color\magenta\heartsuit\] (btw you look pretty cool to me ;) )

OpenStudy (anonymous):

Haha thank you so much. :) I'm horrible with providing examples so thanks. (and that's the joke really, since nobbody ever agrees that im not cool, apparently)